OpenStudy (anonymous):

0.37 L = _____mL (1 point).037 3.7 37 370

OpenStudy (anonymous):

@iNeedTheHelp @jim_thompson5910 @Hero

OpenStudy (anonymous):

@Hero help me please!

jimthompson5910 (jim_thompson5910):

1 L = 1000 mL 0.37*1 L = 0.37*1000 mL 0.37 L = ??? mL

OpenStudy (anonymous):

whats the answer?

jimthompson5910 (jim_thompson5910):

multiply 0.37 and 1000 to get your answer

OpenStudy (anonymous):

thank you!!!